OpenStudy (anonymous):

Part A: Explain why the x-coordinates of the points where the graphs of the equations y = 8x and y = 2x + 2 intersect are the solutions of the equation 8x = 2x + 2. (4 points) Part B: Make tables to find the solution to 8x = 2x + 2. Take the integer values of x between −3 and 3. (4 points) Part C: How can you solve the equation 8x = 2x + 2 graphically? (2 points)

OpenStudy (anonymous):

@Qwertty123 can you help with this one?

OpenStudy (jameshorton):

if they y values are the same then you can equate these 2 equations y = 8x and y = 2x + 2 8x=2x+2

OpenStudy (qwertty123):

The first graph is the set of all points where y = 8x. The second graph is the set of points where y = 2x + 2. If a point is on both graphs, then both equations will be true at the same time. So we will have 8x = y and y = 2x + 2, so 8x = 2x + 2. @jameshorton said above! :D

OpenStudy (qwertty123):

(B) x . 8x . 2x + 2 -3 . -24 . -4 -2 . -16 . -2 -1 . -8 . . 0 0 . . 0 . . 2 1 . . 8 . . 4 2 . 16 . . 6 3 . 24 . . 8 From this you can see that the solution is between 0 and 1. Using algebra we can solve it easily: 8x = 2x + 2 => 6x = 2 => x = 1/3.

OpenStudy (anonymous):

wow you are fast!

OpenStudy (qwertty123):

Part C? Do you know what you have to do? (C) Graph the two lines from part A and find the x-coordinate of the intersection.

OpenStudy (anonymous):

First i have a question about Part b and for part c i dont have any graphs to do this on.

OpenStudy (anonymous):

what do the dot mean in part b?

OpenStudy (anonymous):

@Qwertty123

OpenStudy (qwertty123):

graph 1 => y = 8^x graph 2 => y = 2^x +2 plug in different x's between -3 and 3 .... and find different y's corresponding to those x's... when the y's become equal ... it will be the required solution .... that x is the solution.

OpenStudy (anonymous):

is this part b or part c

OpenStudy (anonymous):

okay so that was the table then?

OpenStudy (qwertty123):

LOL I am getting confused! XD @welshfella

OpenStudy (anonymous):

me too

OpenStudy (qwertty123):

@tHe_FiZiCx99

OpenStudy (welshfella):

i dont see how you are confused @Qwerty123 You've answered the question perfectly

OpenStudy (qwertty123):

I did? Oh....lol...To be honest my brain is not awake yet...